# EchoONE: Segmenting Multiple echocardiography Planes in One Model

### v2

@@ Abstract
In clinical practice of echocardiography examinations, multiple planes containing the heart structures of different view are usually required in screening, diagnosis and treatment of cardiac disease. AI models for echocardiography have to be tailored for each specific plane due to the dramatic structure differences, thus resulting in repetition development and extra complexity. Effective solution for such a multi-plane segmentation (MPS) problem is highly demanded for medical images, yet has not been well investigated. In this paper, we propose a novel solution, EchoONE, for this problem with a SAM-based segmentation architecture, a prior-composable mask learning (PC-Mask) module for semantic-aware dense prompt generation, and a learnable CNN-branch with a simple yet effective local feature fusion and adaption (LFFA) module for SAM adapting. We extensively evaluated our method on multiple internal and external echocardiography datasets, and achieved consistently state-of-the-art performance for multi-source datasets with different heart planes. This is the first time that the MPS problem is solved in one model for echocardiography data. The code will be available at https://github.com/a2502503/EchoONE.
@@ 1 Introduction
Figure 1: Comparison of existing solutions for the Multi-plane segmentation problem in echocardiogprahic images. The performance of classic U-Net significantly decreases when trained on multiple planes together (b), compared to training on each plane individually (a). The direct application of SAM on different echocardiographic planes yields poor results (c). EchoONE demonstrates excellent segmentation performance across all planes (d).
Learning-based segmentation methods have been well developed for different medical images to help assist disease screening, diagnosing, and treatment [33, 31]. It’s usually the case that an organ is scanned from multiple planes of different orientations. For example, in echocardiography examination, various long axis and short axis planes were required to conduct a reliable evaluation of the heart structure and its functionality [27]. Segmentation models for echocardiography have to be tailored for each specific plane due to the dramatic structure differences, thus resulting in repetition development and extra complexity. A uniform model for multiple plane segmentation (MPS) is highly demanded to easy the development and deployment. However, this problem has not been well investigated.
Existing solutions for the MPS problem include multi-branch architecture in model design and combination of multiple plane data for training. The multi-branch architecture usually leverage some cross-view attention/consistency [10, 34] to help improve the performance. Each plane is still treated separately, and is difficulty to generalize to other new planes. When multiple planes are combined to train a uniform segmentation model, performance usually drop obviously when comparing the plane-tailored model. As shown in Row 1 and Row 2 of Figure 1 the results for segmentation of four different heart planes in echocardiography. The model trained with all planes together obtain clearly inferior performance to the plane-specific model.
Segment Anything Model (SAM) [17], which generalizes well for natural images and maybe a potential solution for the MPS problem. However, it fails for medical images due to the large semantic gap. Row 3 of Figure 1 demonstrated the results of SAM when tested with images of the four echocardiographic planes. Without fine-tuning, SAM fail for all these planes. Existing SAM-derived methods in medical images focus on adapting/fine-tuning to a specific scenario [3, 39] or proposing new prompting generation ways [35, 13]. The MPS problem has not been focused in these efforts despite its potential.
@@
    We introduce a local feature fusion and adaption (LFFA) module for feature interaction and fusion with the image encoder and the mask decoder, which help improving segmentation and accelerating convergence.
@@
The results of internal and external validation using three public datasets, and four private datasets demonstrates that our EchoONE can achieve consistently state-of-the-art segmentation performance for left ventricle, left atrium, and myocardium from six echocardiographic planes.
@@ 2.1 Multi-plane segmentation
Multi-plane analysis of medical images enables a comprehensive evaluation of the organ of interest in clinical practice, therefore raising the requirement for multi-plane segmentation (MPS) of medical images. Existing solutions can be attributed into two categories: multi-branch architecture with each branch dealing with one plane, and a uniform model that is trained using combination of multi-plane image-mask pairs. While the later often leads to performance drop, efforts have been devoted into the former one. [9] designed a network with three segmentation branches for each of the sagittal, coronal, and axial planes for brain tumor segmentation. Then fused 3D results of the three planes were then employed to improve the segmentation of each plane in turn. [43] presents multi-plane representation learning for volume medical data by exploring plane discrepancy and dependency within three parallel branches for sagittal, coronal, and axial views, resulting improved 3D segmentation. TransFusion [22] introduced multiscale cross-view context modeling and dependency mining between two parallel branches for the segmentation of short and long axis cardiac MRI images.
@@
Figure 2: Overview of our EchoONE, which contains a SAM-based segmentation architecture (middle), a prior-composable mask learning (PC-Mask) module (top) for dense prompt generation, and a learnable CNN-based local feature branch (bottom) for SAM encoder tunning and decoder adapting.
@@ 2.2 SAM in medical image tasks
SAM is a robust zero-shot segmentation foundation model that has demonstrated exceptional performance across various natural image tasks [41, 36, 28]. However, directly applying SAM to medical imaging often results in a significant loss of generalization capability [14, 25, 4]. This is largely due to the fact that, in contrast to natural images, medical images typically contain low contrast structures, high level noises, indistinct tissue boundaries, thus leading to inherent disparities in data distribution. Currently, many studies focus on fine-tuning SAM to restore its robust segmentation capability in medical imaging tasks. MedSAM [23] constructed a large-scale medical image dataset and utilized bounding box prompts to guide SAM. AutoSAM [32] replaces the interactive input prompts with a learnable encoder, enabling a fully automated procedure. SAMed [42] applies a low-rank-based (LoRA) fine-tuning strategy to the image encoder and uses the default prompt embedding to achieve prompt-free results. SAMMed-2D [7] and MSA [37] both employ Adapter blocks to fine-tune SAM. SAMMed-2D adds an Adapter Layer to each transformer block of the image encoder, adjusting the feature across both channel and spatial dimensions. MSA integrates prompt information into the adaptation process. SAM-LST [2] is inspired by another fine-tuning method, Ladder Side-Tuning (LST), and introduces an additional CNN encoder, providing a new approach for combining CNN with SAM. SAMUS [21] employs both LST and Adapter Tuning methods, using position and feature adapters to transition SAM to medical domains.
In summary, these SAM-based methods facilitate the transition of SAM from natural images to medical images, demonstrating strong feature extraction and generalization ability. However, none have been designed for the MPS problem, especially for the multiple echocardiography planes, where significant structural differences exist between planes and the heavy noise and low contrast may deteriorate the prompting and the mask decoder in SAM.
@@ 2.3 Prompt engineering in SAM
The configuration of prompts significantly impacts the effectiveness of SAM models [14], and has become the focus of recent efforts. SAMAug [8] generates enhanced point prompts that provide more comprehensive user intent information. For remote sensing instance segmentation, RSPrompter [5] introduces a class-based prompt generator, enabling the output mask aligned with the specific global semantics. PA-SAM [38] extracts information from the image for a prompt adapter, generating additional sparse and dense prompts to optimize the decoder. In the medical imaging domain, MSA [37] constructs an adapter module focused on prompt integration, while ProMISe [19] experiments with combining point coordinates and image embeddings to refine prompt information. SurgicalSAM [40] integrates surgical instrument-specific information with SAM’s pretrained knowledge via a learnable prototype-based class prompt module, achieving efficient, lightweight tuning.
These methods add information from the user or the input images to implicitly make a enriched prompt, thus to improve the final performance. We explicitly employ prior semantic knowledge of different planes to sense the structure of the input image and generate the mask prompt.
@@ 3.1 Overview
Figure 2 shows the overview architecture of EchoONE. EchoONE mainly consists of three components, a SAM-based segmentation architecture, a prior-composable mask learning (PC-Mask) module for dense prompt generation, and a CNN-based local feature branch for SAM tunning and adapting.
@@ 3.2 Prior-composable mask learning for dense prompt generation
With these mask centers as the structure prior, we aim to customize them for each new image so that the resulted mask is aware of the structure in that image without knowing its plane information. Specifically, for a input image $I_{j}$, its location in the latent space can be represented by its similarity (or distance) to these prototypes $u_{i}$. In our work, the cosine similarity is used:
@@
With the custmoized prior embedding $PE_{j}$ as input, we design a light-weight U-Net to learn the mask of the input image using a combination of dice loss and binary cross-entropy Loss $L_{PCM}$. The output $PCM_{j}$ is used as our dense prompt for SAM.
@@
It’s worth to note that our PC-Mask is not derived directly from the input image, therefore the semantics differs from traditional segmentation network and could be an effective guidance for SAM. Besides, no plane information is used during the generation, which makes our model capable of dealing the MPS problem in one uniform model.
@@ 3.3 CNN branch for tuning and adaption
Figure 4: Local feature fusion and adaption (LFFA). The output of each CNN layer (local feature) and image embededings from a previous transformer is first concatenated and then mixed with a convolution layer to obtain the fused image embedding.
To fully leverage the capability of SAM without resourcing-intensive retraining, an auxiliary branch for tunning and adapting SAM to a new scenario is necessary. Inspired by previous SAM-based work [21, 2], we design a learnable CNN branch as shown in the bottom of Figure 4 for this. Our CNN branch is constituted of three parts: the Residual block for local feature extraction, the CNN block with cross-branch attention for tuning of the image encoder, and the transform blocks to fusion the local feature to adapt the mask decoder to our task.
Figure 5: (a) Remapping annotations to unified mask representation. (b) Preprocessings to generate LV cavity from the annotation of MYO in mulitple planes. Landmarks are quired to be detected first for planes of 2CH, 3CH and 4CH.
@@
where $\mathbf{DTrans}$ is the transformer block in the mask decoder of SAM. Here the other two inputs of the transformer block, i.e., the position embedding and the point prompt are ignored for simplicity. We have observed in our experiments that such a LFFA design not only integrates local feature better, but also accelerates the model’s convergence.
@@ D $1774^{(2,3)}$ $210^{(2,3)}$ $1758^{(2,3)}$ $1387^{(2,3)}$
Table 1: Summary of the plane and annotation information for three public datasets and four private datasets. The abbreviations A-D represent different datasets. The upper script (1,2,3) are short for annotation of the left ventricle, myocardium, and left atrium.
@@ 4.1 Datasets and evaluation metrics
Dataset details: Details of the plane information and the annotation can be seen in Table 1. There are a total of six echocardiogrpahic planes, i.e., apical 2-chamber (2CH), apical 3-chamber (3CH), apical 4-chamber (4CH), parasternal short axis basal level (PSAX-B), parasternal short axis mid-ventricle level (PSAX-M), parasternal short axis apical level (PSAX-A). Given the fact that the three PSAX planes share similarities, we treat them as one PSAX plane in this work. In the table, the image number of each plane in the dataset and how each plane is annotated are demonstrated. Take CAMUS as example, two planes of 2CH and 4CH are included. For each image, the left ventricle (LV), myocardium (MYO) and the left atrium (LA) are annotated.
Unified mask representation for inconsistent annotation protocols: Since our datasets have different sources, contains different planes, the annotation information differs a lot across datasets. It’s necessary to first unify their annotation representation so that a uniform model can be developed based on them. In this work, we remap all annotations according to the following protocols: 0 for background, 1 for LV, 2 for LV cavity, and 3 for MYO. Since all planes show the heart structure from a specific orientation, such as protocol is applicable to all planes, and to all datasets as well. Also see Figure 5 (a) for examples of this annotation remapping. Besides, for those image with only MYO annotated, we generate the masks for LV cavity by detection landmarks and cavity filling (as shown in Figure 5 (b)).
@@
Training and evaluation: The datasets are used as follows during training and evaluation. Datasets CAMUS, A, B, C, and D are used as internal data, where images of 80% subjects for each are randomly selected for model training, 10% for validation during training, and the rest 10% are used as internal evaluation. Datasets HMC_QU and EchoNet-Dynamic are only used for external evaluation to test the generalization of our model.
@@ 4.2 Implementation details
Before training, the parts inherited from SAM were initialized with the pre-trained weights of SAM on the SA-1B dataset using a ViT-B backbone, while the remaining parameters were initialized randomly. During the training process, we employed a combination of random rotation, scaling, contrast adjustment and gamma enhancement for data augmentation, each with a probability of 0.5. All images resolution were resized to 256×256. We trained our datasets for 100 epochs. The Adam optimizer is used, with $\beta_{1}$, $\beta_{2}$ set to 0.9 and 0.999, and an initial learning rate of $41e-$ 4. All our implementation is in PyTorch and we train our models on single NVIDIA 48G A6000 GPU.
@@
4.3 Results
We conduct extensive experiments using internal and external datasets to demonstrate the effectiveness of our method for the MPS problem. Existing methods used for comparison include four traditional models: CNN-based U-Net [30] and DeepLabV3+ [6], transformer-based SwinUNet [1], and the CNN-transformer hybrid model H2Former [11], as while as six SAM-based methods: the original SAM [17], MedSAM [23], SAMed [42], SAM-Med2D [7], MSA [37], and SAMUS [21]. Except SAM, which is applied directly to the test data, the rest methods are trained using the same datasets as those for training EchoONE.

### v3

@@ Abstract
In clinical practice of echocardiography examinations, multiple planes containing the heart structures of different view are usually required in screening, diagnosis and treatment of cardiac disease. AI models for echocardiography have to be tailored for each specific plane due to the dramatic structure differences, thus resulting in repetition development and extra complexity. Effective solution for such a multi-plane segmentation (MPS) problem is highly demanded for medical images, yet has not been well investigated. In this paper, we propose a novel solution, EchoONE, for this problem with an SAM-based segmentation architecture, a prior-composable mask learning (PC-Mask) module for semantic-aware dense prompt generation, and a learnable CNN-branch with a simple yet effective local feature fusion and adaption (LFFA) module for SAM adapting. We extensively evaluated our method on multiple internal and external echocardiography datasets and achieved consistently state-of-the-art performance for multi-source datasets with different heart planes. This is the first time the MPS problem has been solved in one model for echocardiography data. The code will be available at https://github.com/a2502503/EchoONE.
@@ 1 Introduction
Learning-based segmentation methods have been well developed for different medical images to help assist disease screening, diagnosing, and treatment [35, 33]. It’s usually the case that an organ is scanned from multiple planes of different orientations. For example, in the echocardiography examination, various long- and short-axis planes were required to perform a reliable evaluation of the structure and functionality of the heart [28]. Segmentation models for echocardiography have to be tailored for each specific plane due to dramatic structure differences, thus resulting in repetition development and extra complexity. A uniform model for multiple plane segmentation (MPS) is highly demanded to easy the development and deployment. However, this problem has not been well investigated.
Figure 1: Comparison of existing solutions for the Multi-plane segmentation problem in echocardiographic images. The performance of classic U-Net significantly decreases when trained on multiple planes together (b), compared to training on each plane individually (a). The direct application of SAM on different echocardiographic planes yields poor results (c). EchoONE demonstrates excellent segmentation performance across all planes (d).
Existing solutions for the MPS problem include multi-branch architecture in model design and combination of multiple plane data for training. The multi-branch architecture usually leverage some cross-view attention/consistency [10, 36] to help improve the performance. Each plane is still treated separately, and is difficult to generalize to other new planes. When multiple planes are combined to train a uniform segmentation model, performance usually drops obviously when comparing the plane-tailored model. As shown in Row 1 and Row 2 of Figure 1 the results for segmentation of four different heart planes in echocardiography. The model trained with all planes together obtains clearly inferior performance to the plane-specific model.
Segment Anything Model (SAM) [17], which generalizes well for natural images and may be a potential solution for the MPS problem. However, it fails for medical images due to the large semantic gap. Row 3 of Figure 1 demonstrated the results of SAM when tested with images of the four echocardiographic planes. Without fine-tuning, SAM fails for all these planes. Existing SAM-derived methods in medical images focus on adapting/fine-tuning to a specific scenario [3, 41] or proposing new prompting generation ways [37, 13]. The MPS problem has not been focused in these efforts despite its potential.
@@
    We introduce a local feature fusion and adaption (LFFA) module for feature interaction and fusion with the image encoder and the mask decoder, which helps improve segmentation and accelerate convergence.
@@
The results of internal and external validation using three public datasets, and four private datasets demonstrate that our EchoONE can achieve consistently state-of-the-art segmentation performance for left ventricle, left atrium, and myocardium from six echocardiographic planes.
@@ 2.1 Multi-plane segmentation
Multi-plane analysis of medical images enables a comprehensive evaluation of the organ of interest in clinical practice, therefore raising the requirement for multi-plane segmentation (MPS) of medical images. It is noteworthy that the challenge of multi-plane analysis transcends medical imaging domains. [27] integrates a semantic neural radiance field in conjunction with a video instance segmentation model to segment the appearance of the same scene from multiple view directions. [30] uses three encoders with a shared latent space for multi-view radar segmentation. Existing solutions in the domain of medical image segmentation can be attributed into two categories: multi-branch architecture with each branch dealing with one plane, and a uniform model that is trained using combination of multi-plane image-mask pairs. While the latter often leads to performance drop, efforts have been devoted to the former one. [9] designed a network with three segmentation branches for each of the sagittal, coronal, and axial planes for brain tumor segmentation. Then fused 3D results of the three planes were then employed to improve the segmentation of each plane in turn. [45] presents multi-plane representation learning for volume medical data by exploring plane discrepancy and dependency within three parallel branches for sagittal, coronal, and axial views, resulting in improved 3D segmentation. TransFusion [22] introduced multiscale cross-view context modeling and dependency mining between two parallel branches for the segmentation of short and long axis cardiac MRI images.
@@
Figure 2: Overview of our EchoONE, which contains a SAM-based segmentation architecture (middle), a prior-composable mask learning (PC-Mask) module (top) for dense prompt generation, and a learnable CNN-based local feature branch (bottom) for SAM encoder tuning and decoder adapting.
@@ 2.2 SAM in medical image tasks
SAM is a robust zero-shot segmentation foundation model that has demonstrated exceptional performance on various natural image tasks [43, 38, 29]. However, directly applying SAM to medical imaging often results in a significant loss of generalization capability [14, 25, 4]. This is largely due to the fact that, in contrast to natural images, medical images typically contain low contrast structures, high level noises, indistinct tissue boundaries, thus leading to inherent disparities in data distribution. Currently, many studies focus on fine-tuning SAM to restore its robust segmentation capability in medical imaging tasks. MedSAM [23] constructed a large-scale medical image dataset and utilized bounding box prompts to guide SAM. AutoSAM [34] replaces the interactive input prompts with a learnable encoder, enabling a fully automated procedure. SAMed [44] applies a low-rank-based (LoRA) fine-tuning strategy to the image encoder and uses the default prompt embedding to achieve prompt-free results. SAMMed-2D [7] and MSA [39] both employ Adapter blocks to fine-tune SAM. SAMMed-2D adds an Adapter Layer to each transformer block of the image encoder, adjusting the feature across both channel and spatial dimensions. MSA integrates prompt information into the adaptation process. SAM-LST [2] is inspired by another fine-tuning method, Ladder Side-Tuning (LST), and introduces an additional CNN encoder, providing a new approach for combining CNN with SAM. SAMUS [21] employs both LST and Adapter Tuning methods, using position and feature adapters to transition SAM to medical domains.
In summary, these SAM-based methods facilitate the transition of SAM from natural images to medical images, demonstrating strong feature extraction and generalizability. However, none have been designed for the MPS problem, especially for the multiple echocardiography planes, where significant structural differences exist between planes and the heavy noise and low contrast may deteriorate the prompting and the mask decoder in SAM.
@@ 2.3 Prompt engineering in SAM
The configuration of the prompts significantly impacts the effectiveness of SAM models [14], and has become the focus of recent efforts. SAMAug [8] generates enhanced point prompts that provide more comprehensive user intent information. For remote sensing instance segmentation, RSPrompter [5] introduces a class-based prompt generator, enabling the output mask to align with the specific global semantics. PA-SAM [40] extracts information from the image for a prompt adapter, generating additional sparse and dense prompts to optimize the decoder. In the medical imaging domain, MSA [39] constructs an adapter module focused on prompt integration, while ProMISe [19] experiments with combining point coordinates and image embeddings to refine prompt information. SurgicalSAM [42] integrates surgical instrument-specific information with SAM’s pre-trained knowledge via a learnable prototype-based class prompt module, achieving efficient and lightweight tuning.
These methods add information from the user or the input images to implicitly make a enriched prompt, thus improving the final performance. We explicitly employ prior semantic knowledge of different planes to sense the structure of the input image and generate the mask prompt.
@@ 3.1 Overview
Figure 2 shows the overview architecture of EchoONE. EchoONE mainly consists of three components, a SAM-based segmentation architecture, a prior-composable mask learning (PC-Mask) module for dense prompt generation, and a CNN-based local feature branch for SAM tuning and adapting.
@@ 3.2 Prior-composable mask learning for dense prompt generation
With these mask centers as the structure prior, we aim to customize them for each new image so that the resulted mask is aware of the structure in that image without knowing its plane information. Specifically, for an input image $I_{j}$, its location in the latent space can be represented by its similarity (or distance) to these prototypes $u_{i}$. In our work, the cosine similarity is used:
@@
With the customized prior embedding $PE_{j}$ as input, we design a light-weight U-Net to learn the mask of the input image using a combination of dice loss and binary cross-entropy Loss $L_{PCM}$. The output $PCM_{j}$ is used as our dense prompt for SAM.
@@
It’s worth to note that our PC-Mask is not derived directly from the input image, therefore the semantics differs from traditional segmentation network and could be an effective guidance for SAM. Besides, no plane information is used during the generation, which makes our model capable of dealing with the MPS problem in one uniform model.
@@ 3.3 CNN branch for tuning and adaption
Figure 4: Local feature fusion and adaption (LFFA). The output of each CNN layer (local feature) and image embeddings from a previous transformer is first concatenated and then mixed with a convolution layer to obtain the fused image embedding.
To fully leverage the capability of SAM without resourcing-intensive retraining, an auxiliary branch for tuning and adapting SAM to a new scenario is necessary. Inspired by previous SAM-based work [21, 2], we design a learnable CNN branch as shown in the bottom of Figure 4 for this. Our CNN branch is constituted of three parts: the Residual block for local feature extraction, the CNN block with cross-branch attention for tuning of the image encoder, and the transform blocks to fusion the local feature to adapt the mask decoder to our task.
@@
where $\mathbf{DTrans}$ is the transformer block in the mask decoder of SAM. Here the other two inputs of the transformer block, i.e., the position embedding and the point prompt, are ignored for simplicity. We have observed in our experiments that such an LFFA design not only integrates local feature better, but also accelerates the model’s convergence.
Figure 5: (a) Remapping annotations to unified mask representation. (b) Preprocessing to generate LV cavity from the annotation of MYO in multiple planes. Landmarks are required to be detected first for planes of 2CH, 3CH and 4CH.
@@ D $1774^{(2,3)}$ $210^{(2,3)}$ $1758^{(2,3)}$ $1387^{(2,3)}$
Table 1: Summary of the plane and annotation information for three public datasets and four private datasets. The abbreviations A-D represent different datasets. The upper script (1,2,3) are short for annotation of the left atrium, left ventricle and myocardium.
@@ 4.1 Datasets and evaluation metrics
Dataset details: Details of the plane information and annotation can be seen in Table 1. There are a total of six echocardiographic planes, i.e., apical 2-chamber (2CH), apical 3-chamber (3CH), apical 4-chamber (4CH), parasternal short axis basal level (PSAX-B), parasternal short axis mid-ventricle level (PSAX-M), parasternal short axis apical level (PSAX-A). Given the fact that the three PSAX planes share similarities, we treat them as one PSAX plane in this work. In the table, the image number of each plane in the dataset and how each plane is annotated are demonstrated. Taking CAMUS as an example, two planes of 2CH and 4CH are included. For each image, the left ventricle (LV), myocardium (MYO) and the left atrium (LA) are annotated.
Unified mask representation for inconsistent annotation protocols: Since our datasets have different sources, contain different planes, the annotation information differs a lot across datasets. It’s necessary to first unify their annotation representation so that a uniform model can be developed based on them. In this work, we remap all annotations according to the following protocols: 0 for background, 1 for LV, 2 for LV cavity, and 3 for MYO. Since all planes show the heart structure from a specific orientation, such as protocol is applicable to all planes, and to all datasets as well. Also see Figure 5 (a) for examples of this annotation remapping. Besides, for those image with only MYO annotated, we generate the masks for LV cavity by detecting the landmarks and filling the cavity (as shown in Figure 5 (b)).
@@
Training and evaluation: The datasets are used as follows during training and evaluation. Datasets of CAMUS, A, B, C, and D are used as internal data, where images of 80% subjects for each are randomly selected for model training, 10% for validation during training, and the rest 10% are used as internal evaluation. Datasets HMC_QU and EchoNet-Dynamic are only used for external evaluation to test the generalization of our model.
@@ 4.2 Implementation details
Before training, the parts inherited from SAM were initialized with the pre-trained weights of SAM on the SA-1B dataset using a ViT-B backbone, while the remaining parameters were initialized randomly. During the training process, we employed a combination of random rotation, scaling, contrast adjustment, and gamma enhancement for data augmentation, each with a probability of 0.5. All images were resized to 256×256. We trained our datasets for 100 epochs. The Adam optimizer is used, with $\beta_{1}$, $\beta_{2}$ set to 0.9 and 0.999, and an initial learning rate of $41e-$ 4. All our implementation is in PyTorch and we train our models on single NVIDIA 48G A6000 GPU.
@@ 4.3 Results
We conducted extensive experiments using internal and external datasets to demonstrate the effectiveness of our method for the MPS problem. Existing methods used for comparison include four traditional models: CNN-based U-Net [32] and DeepLabV3+ [6], transformer-based SwinUNet [1], and the CNN-transformer hybrid model H2Former [11], as while as six SAM-based methods: the original SAM [17], MedSAM [23], SAMed [44], SAM-Med2D [7], MSA [39], and SAMUS [21]. Except SAM, which is applied directly to the test data, the rest of the methods are trained using the same datasets as those used to train EchoONE.